# LinkAlign: Scalable Schema Linking for Real-World Large-Scale Multi-Database Text-to-SQL

### v2

@@ Abstract
Schema linking is a critical bottleneck in achieving human-level performance in Text-to-SQL tasks, particularly in real-world large-scale multi-database scenarios. Addressing schema linking faces two major challenges: (1) Database Retrieval: selecting the correct database from a large schema pool in multi-database settings, while filtering out irrelevant ones. (2) Schema Item Grounding: accurately identifying the relevant tables and columns from within a large and redundant schema for SQL generation. To address this, we introduce LinkAlign, a novel framework that can effectively adapt existing baselines to real-world environments by systematically addressing schema linking. Our framework comprises three key steps: multi-round semantic enhanced retrieval and irrelevant information isolation for Challenge 1, and schema extraction enhancement for Challenge 2. We evaluate our method performance of schema linking on the SPIDER and BIRD benchmarks, and the ability to adapt existing Text-to-SQL models to real-world environments on the SPIDER 2.0-lite benchmark. Experiments show that LinkAlign outperforms existing baselines in multi-database settings, demonstrating its effectiveness and robustness. On the other hand, our method ranks highest among models excluding those using long chain-of-thought reasoning LLMs. This work bridges the gap between current research and real-world scenarios, providing a practical solution for robust and scalable schema linking. The codes are available at https://github.com/Satissss/LinkAlign.
@@ 1 Introduction
Figure 1: Comparision across methods on Spider 2.0-lite benchmark. Our method ranks highest, excluding those using long CoT reasoning LLMs, by simply applying LinkAlign to the DIN-SQL base model.
@@
Text-to-SQL Zhong et al. (2017); Wang et al. (2017); Cai et al. (2017); Qin et al. (2022) aims to enable non-expert users to retrieve data effortlessly by automatically translating natural language questions into accurate SQL queries. Recent advances in large language models (LLMs) have led to notable improvements in Text-to-SQL benchmarks Sun et al. (2023); Pourreza et al. (2024), showcasing their growing capabilities in understanding and generating SQL queries. However, despite these advances, existing methods often struggle with real-world applications due to enterprise databases often containing massive redundant schemas and multi-database settings (e.g., local SQLite and cloud BigQuery). It faces significant failures in adapting existing Text-to-SQL models to large-scale multi-database scenarios, largely due to schema linking, i.e., identifying the necessary database schemas (tables and columns) from large volumes of database schemas for a user query Wang et al. (2019); Guo et al. (2019); Talaei et al. (2024). The underlying reasons for these failures remain unexplored, leaving a gap in addressing the real-world Text-to-SQL tasks.
@@
To understand the failures, we perform a comprehensive error analysis and conclude that schema linking failures stem from two major challenges. Challenge 1 - Database Retrieval : how to select the correct database from a large schema pool in multi-database settings, while filtering out irrelevant ones. Existing researches often overlook this challenge as they always assume that single-database schemas are small-scale and can be directly fed into models for efficient processing. Challenge 2 - Schema Item Grounding : how to accurately identify the relevant tables and columns from within a large and redundant schema for SQL generation. The post-retrieval phase must handle complex database schemas, where the large volume of semantically similar tables and columns increases the risk of overlooking critical schema items necessary for generating accurate SQL queries.
@@
Motivated by these factors, we propose LinkAlign, a novel framework that systematically addresses the challenges of schema linking in large-scale, multi-database scenarios. Specifically, our framework consists of three major steps. To address Challenge 1, our approach focuses first on (1) Retrieving Potential Database Schemas through multi-round semantic alignment by query rewriting. This step infers missing schema elements based on the retrieved result evaluation and rewrites the query to align with the missing database schema. Then our approach centers on (2) isolating irrelevant schema information to reduce noise by response filtering. This step filters the most appropriate database from a set of candidates, minimizing interference from irrelevant schemas, and streamlining the downstream processing pipeline. To address Challenge 2, our approach directs efforts towards (3) extracting schemas for SQL generation through identifying tables and columns by schema parsing. This step pushes the boundary of schema linking in large-scale databases by introducing advanced in-context learning techniques and carefully designed prompts for LLMs. To balance efficiency and accuracy, we introduce two distinct implementation strategies for our framework, i.e., pipeline mode and iterative mode. For the pipeline mode, a single LLM call is used to process each step, which is particularly well suited for scenarios where efficiency is critical. In contrast, the iterative mode adopts a multi-round multi-agent collaboration, which prioritizes accuracy and excels at handling complex and ambiguous queries, making it ideal for challenging Text-to-SQL tasks.
To better evaluate the model’s capabilities of schema linking, we construct AmbiDB, a variant of the Spider Yu et al. (2018) benchmark, designed to simulate the ambiguity and complexity of Text-to-Sql in real-world environment. We evaluate the schema linking capability of our approach on the SPIDER, BIRD Li et al. (2024a) benchmarks and the AmbiDB datasets proposed by ourselves, comparing it with existing baseline methods. Experimental results show that our approach achieves state-of-the-art performance across all metrics under multi-database settings. Additionally, we apply LinkAlign to the classic DIN-SQL method and conduct experiments on the SPIDER 2.0 Lei et al. (2024) benchmark to validate the effectiveness in adapting existing baselines to large-scale, multi-database scenarios. The results in Figure 1 show that LinkAlign ranks highest excluding those using long CoT reasoning LLMs.
2 Related Work
@@ 3 Error Analysis
To better understand the gap between existing research and real-world environment, we test 500 samples from the Spider dataset and analyze common error types in existing baseline methods for Text-to-SQL tasks. In particular, the model needs to handle schemas from all databases, rather than small-scale schemas from a single database, which is more aligned with real-world scenarios. The results show that schema linking errors are the main cause of Text-to-SQL failures, with an error rate greater than 60%. To better understand the failures, we manually examined the erroneous samples and attributed the four common error types to two major challenges proposed in the Introduction. Appendix A provides more details of the results.
@@
Error 2: Referring Irrelevant Databases (Database Retrieval). Unlike Error 1 that focuses on the missing target database, Error 2 centers on the irrelevant database noise introduced by imprecise retrieval under multi-database settings. This error indicates that the model refers to incorrect schemas from unrelated databases when generating SQL, accounting for 13.3% of the failures. For example, the user intends to query "The first name of students who have both cats and dogs". However,the generated SQL incorrectly infers $People.first\_name$ from an unrelated database while overlooking the ground truth $Student.f\_name$, even though both are successfully retrieved and fed into the model. Without isolating irrelevant database information, the model tends to mistakenly select seemingly more appropriate schemas from unrelated databases, leading to SQL execution failures.
@@ 4 Methodology
This section presents the LinkAlign framework for schema linking and the implementation strategies used to realize it. Our framework consists of three major steps designed to address the challenges in large-scale, multi-database scenarios. The implementation strategies provide two complementary approaches to execute these steps balancing efficiency and accuracy.
@@ 4.1 Background
4.2 LinkAlign Framework
The LinkAlign framework can effectively adapt existing models to real-world environments by systematically addressing the challenges of schema linking. The framework consists of three main steps that facilitate the downscaling of database schemas and enhance schema linking capabilities.
Step one: retrieve potential database schemas. To mitigate the Error 1, we propose a method for multi-round semantically enhanced retrieval. In this step, LLMs are used to evaluate the retrieval results and rewrite the input query based on the inferred missing database schemas. Dynamically adjusting the retrieval strategy to align semantically with missing database schemas enables high retrieval performance with fewer iterations. This approach is particularly effective in scenarios with limited retrieval batches and restricted scale. {newshaded}User Query $Q_{0}$: Which semester the master and the bachelor both got enrolled in?
@@ 4.2 LinkAlign Framework
[2] Rewrite $Q_{2}$: In a database with enrollment_records, how to find semesters where both master and bachelor students enrolled? Group by semester and filter for overlapping enrollments. The rewritten queries are embedded into vectors and then used to retrieve the relevant database schemas. The retrieval results are sorted and summarized based on the number of rewrites and similarity scores:
where $T$ represents the number of query rewrites. By rewriting queries and enhancing their semantic representation, this approach improves the alignment between user queries and database schemas, ensuring more accurate retrieval outcomes.
Step two: isolate irrelevant schema information. To mitigate Error 2, we attempt to filter out irrelevant schemas from the retrieved multi-database results. Once the retrieved results $Z$ containing schemas from irrelevant databases, the next step is to precisely locate the target database $D_{t}$ while filtering out irrelevant ones. This component is designed to minimize interference from extraneous schemas, which can obscure the generation of accurate SQL queries. To achieve this, the framework evaluates the relevance of each database $D_{i}$ from retrieved results $Z$ based on the original user query $Q_{0}$. The most likely target database $D_{t}$ is selected based on its analyzed alignment with the query.
@@
where M denotes the LLMs used for analysis. By focusing on the most appropriate database, this step isolates unrelated schemas information from being included in subsequent operations. This filtering is essential, as irrelevant schemas can unnecessarily extend the input context, increasing computational costs and diminishing SQL generation performance. By selecting only the target database, the framework ensures that schema linking focuses on relevant information, thereby improving both accuracy and efficiency. Additionally, our approach remains effective in single-database settings by adjusting the objective to filter out irrelevant database schemas rather than those from unrelated databases, as explained in Appendix C.
Step three: extract schemas for SQL generation. To mitigate the Error 3 and 4, we need to identify the correct tables and columns for SQL generation. This step mimics the manual process of database schema extraction, but it can be automated using reasoning techniques such as Chain-of-thought. Specifically, given filtered database schema $S_{\hat{u}}$ from previous steps, we need to identify a subset $S_{\hat{u}}^{\prime}$ that contains the most relevant tables $T_{i}^{\hat{u}}$ and columns $C_{i}^{\hat{u}}$ These elements are selected based on their alignment with the query, ensuring that the resulting schema is both precise and comprehensive.
@@ 4.3 Component Implementation Optimization
Based on the formatted definitions of each step in Section 4.2, we propose two strategies to further implement the key components in each step, with positions shown in the dashed box in Figure 2. These strategies share modular interfaces and the specific calls tailored to different levels of query complexity and database scale. The first one is Single-Prompt Pipeline, which simplifies the implementation of the LinkAlign framework through a single call to LLMs. This approach is particularly well-suited for large-scale scenarios where efficiency is critical. Appendix B provides a detailed explanation of this strategies, while this section focuses on Multi-agent Collaboration strategy, which prioritizes accuracy and provide sufficient capabilities to tackle complex query tasks in real-world environment.
Align Semantics by Query Rewriting. Inspired by the reflection capability of LLMs discussed by Shinn et al., we design an semantic enhanced retrieval approach composed of two step: inferring missing schemas and query rewriting. Specifically, the Schema Auditor first evaluates the retrieval results, infers the missing schemas required for accurate SQL generation (e.g., fields for SELECT/JOIN/WHERE), and concludes with an audit report. Next, the Query Rewriter enriches the question with additional semantic information to better align with ifferred missing database schemas, which ultimately enhances retrieval accuracy.
@@
Reduce Noise by Response Filtering. When multiple candidate databases show minor differences, reaching consensus through multi-agent debates significantly reduces the risk of errors Chan et al. (2023). Inspired by this, we develop a multi-agent debate model composed of Data Analyst and Database Expert. Specifically, the Data Analyst infers the most appropriate database from multiple candidates, and the Database Expert evaluates whether the selection is reasonable and decides retain or not. The debate follows a one-by-one strategy, i.e., starting with the data analyst, after which the two roles present their perspectives in turn. The debate ends when a predefined number of rounds is reached, and then a terminator outputs the consensus database as a final result.
@@
Identify Tables and Columns by Schema Parsing. To enhance schema linking capabilities, we develop a multi-agent debate model based on Simultaneous-Talk-with-Summarizer strategy, i.e., multiple peer participants engage in simultaneous debate, with the final outcomes evaluated by a summarizer. The debate involves two roles: Schema Parser and Data Scientist. In each round, each Schema Parser extracts a clause-specific database schema (e.g. fields for JOIN/WHERE…), and then the Data Scientist verifies all results from Schema Parsers, identifying any omissions or errors. Multi-role participation enhances the recall of tables and columns required for SQL generation, with diverse answers complementing each other to reduce the randomness of single-prompt outputs.

### v3

@@ Abstract
Schema linking is a critical bottleneck in applying existing Text-to-SQL models to real-world, large-scale, multi-database environments. Through error analysis, we identify two major challenges in schema linking: (1) Database Retrieval: accurately selecting the target database from a large schema pool, while effectively filtering out irrelevant ones; and (2) Schema Item Grounding: precisely identifying the relevant tables and columns within complex and often redundant schemas for SQL generation. Based on these, we introduce LinkAlign, a novel framework tailored for large-scale databases with thousands of fields. LinkAlign comprises three key steps: multi-round semantic enhanced retrieval and irrelevant information isolation for Challenge 1, and schema extraction enhancement for Challenge 2. Each stage supports both Agent and Pipeline execution modes, enabling balancing efficiency and performance via modular design. To enable more realistic evaluation, we construct AmbiDB, a synthetic dataset designed to reflect the ambiguity of real-world schema linking. Experiments on widely-used Text-to-SQL benchmarks demonstrate that LinkAlign consistently outperforms existing baselines on all schema linking metrics. Notably, it improves the overall Text-to-SQL pipeline and achieves a new state-of-the-art score of 33.09% on the Spider 2.0-Lite benchmark using only open-source LLMs, ranking first on the leaderboard at the time of submission. The codes are available at https://github.com/Satissss/LinkAlign.
@@ 1 Introduction
Approach Open-source Score (%)
ReFoRCE + o1-preview 30.35
Spider-Agent + Claude-3.7 25.41
Spider-Agent + o3-mini 23.40
DailSQL + GPT-4o 5.68
CHESS + GPT-4o 3.84
DIN-SQL + GPT-4o 1.46
LinkAlign + DeepSeek-R1 33.09
LinkAlign + DeepSeek-V3 24.86
Table 1: Comparision across methods on Spider 2.0-lite benchmark. Our method achieves new SOTA score of 33.09 purely using open-source LLMs.
@@
Text-to-SQL Zhong et al. (2017); Wang et al. (2017); Cai et al. (2017); Qin et al. (2022) aims to enable non-expert users to retrieve data effortlessly by automatically translating natural language questions into accurate SQL queries. Recent advances in large language models (LLMs) have led to notable improvements in Text-to-SQL benchmarks Sun et al. (2023); Pourreza et al. (2024), showcasing their growing capabilities in understanding and generating SQL queries. However, existing methods often fall short in real-world enterprise applications due to difficulties in handling massive redundant schemas and complex multi-database environments (e.g., local and cloud systems). It faces significant failures in adapting existing Text-to-SQL models to large-scale multi-database scenarios largely due to schema linking, i.e., identifying the necessary database schemas (tables and columns) from large volumes of database schemas for user queries Wang et al. (2019); Guo et al. (2019); Talaei et al. (2024). The underlying reasons for these failures remain unexplored, leaving a gap in addressing the real-world Text-to-SQL tasks.
@@
To understand the failures, we conduct a systematic error analysis and identify two major challenges underlying the schema linking. Challenge 1 - Database Retrieval : how to accurately select the target database from a large schema pool, while effectively filtering out irrelevant ones. Existing researches often overlook this challenge as they often assume that schemas from single-database are small-scale and can be directly fed into models for efficient processing. Challenge 2 - Schema Item Grounding : how to precisely identify the relevant tables and columns within complex and redundant schemas for SQL generation. The post-retrieval phase must handle a large volume of semantically similar tables and columns, which increases the risk of overlooking critical items necessary for generating accurate SQL queries.
@@
Motivated by these factors, we propose LinkAlign, a novel framework that systematically addresses the challenges of schema linking in real-world environment. To address Challenge 1, our approach focuses first on (1) Retrieving Potential Database Schemas through multi-round semantic alignment by query rewriting. This step infers missing database schemas from retrieval results by leveraging the LLM’s reflective capabilities, then rewrites the query to align with the ground truth schema semantically. Then our approach centers on (2) isolating irrelevant schema information to reduce noise by response filtering. This step filters out the database noise from a set of candidates, minimizes interference from irrelevant schemas, and streamlines the downstream processing pipeline. To address Challenge 2, our approach directs efforts towards (3) extracting schemas for SQL generation through identifying tables and columns by schema parsing. This step scales schema linking to large-scale databases by introducing advanced reasoning-enhanced prompting techniques like multi-agent debate Chan et al. (2023) and chain-of-thought Wei et al. (2022). To balance efficiency and accuracy, we propose two complementary implementation paradigms: Pipeline and Agent. The pipeline mode executes each step via a process-fixed single LLM call, offering a streamlined, low-latency solution ideal for real-time database queries. In contrast, the agent mode performs multi-turn agent collaboration during inference, harnessing test-time computation to scale schema linking capabilities to databases with massive and complex schemas.
To better evaluate the model’s schema linking capabilities, we automatically construct AmbiDB, a variant of the Spider Yu et al. (2018) benchmark, which introduces a large number of complex synonymous databases to simulate the challenges in large-scale, multi-database scenarios. We perform comprehensive evaluations on the Spider, Bird Li et al. (2024a) and Spider 2.0 Lei et al. (2024) benchmarks. The framework consistently outperforms baselines in all schema linking metrics. By applying LinkAlign to the classic DIN-SQL Pourreza and Rafiei (2024) method, the framework achieves a state-of-the-art score of 33.09 on the Spider 2.0-Lite benchmark using only open-source LLMs, highlighting its effectiveness in tackling schema linking challenges and enhancing the performance of the Text-to-SQL pipeline.
@@ 2 Error Analysis
To better understand the gap between existing research and real-world environments, we evaluate 500 samples from the Spider dataset and analysis common error types when models handling schemas across all databases, rather than limiting the scope to small-scale schemas from a single database. To avoid context-length overflow, we employ vectorized retrieval to extract semantically relevant schemas. The results indicate that schema linking errors are the main cause of Text-to-SQL failures, with an error rate greater than 60%. We manually examined the erroneous samples and identified four error types, which further highlight the two key challenges presented in the Introduction. More details are provided in Appendix B.
@@
Error 2: Referring Irrelevant Databases (Database Retrieval). Unlike Error 1 that focuses on the missing target database, Error 2 centers on the irrelevant schema noise introduced by imprecise retrieval. This error indicates that the model refers to incorrect schemas from unrelated databases when generating SQL, accounting for 13.3% of the failures. For example, the user intends to query "The first name of students who have both cats and dogs". However,the generated SQL incorrectly infers $People.first\_name$ from an unrelated database while overlooking the ground truth $Student.f\_name$, even though both are successfully retrieved and fed into the model. Without isolating irrelevant database information, the model tends to mistakenly select seemingly more appropriate schemas from unrelated databases, leading to SQL execution failures.
@@ 3 Methodology
This section introduces the LinkAlign framework, scaling schema linking to large-scale, multi-database environments through three key steps. The framework begins by (1) retrieving potential database schemas via multi-round semantic alignment through query rewriting, effectively recalling the ground-truth schemas while significantly reducing the candidate pool. Next, (2) isolates irrelevant schema information through response filtering, enabling precise target database localization and noise reduction by discarding unrelated candidates. Finally, it focuses on (3) extracting schemas for SQL generation by identifying necessary tables and columns through schema parsing. To balance efficiency and effectiveness, we provide two complementary implementation modes—Pipeline and Agent—for each step of the framework.
@@ 3.1 Background
3.2 Modular Step Design
This section outlines the modular design of each step, decoupling from implementations to accommodate diverse application scenarios.
Step one: retrieve potential database schemas. To mitigate the exclusion of the ground truth schema (Error 1), we propose a multi-round semantically enhanced retrieval method to recall critical schemas without significantly increasing the retrieval size. This step infers missing schemas from retrieval results by leveraging the LLM’s reflective capabilities, then rewrites the query to align with the ground-truth schema semantically.
Specifically, following each retrieval round, field-level metadata (e.g., type, description, value example) from index nodes are extracted and serialized into structured natural language sequences aligned with LLM processing preferences. The resulting schema representation, combined with the original user query, forms the context, denoted as the tuple ($S_{r_{i}},\ Q_{0}$). Subsequently, we leverage LLMs to evaluate the semantic alignment between the user query and the retrieved schema context, and further infer potentially missing schema elements critical for accurate SQL generation. The inferred schemas are integrated with the original query and optimized to remove redundant or ambiguous expressions. This integration helps reduce hallucination-induced deviations from user intent and improves semantic alignment with the ground-truth schema. The rewritten queries are then embedded into vector representations and used to retrieve relevant database schemas. Finally, the retrieval results are ranked and aggregated based on the number of rewrites and their similarity scores:
where $T$ represents the number of query rewrites. By rewriting queries and enhancing their semantic representation, this approach improves the alignment between user queries and database schemas, ensuring more accurate retrieval outcomes. Dynamically adjusting the retrieval strategy based on feedback enables high retrieval performance with fewer iterations. Concurrently, multi-round iterative optimization enables effective scalability to large-scale databases with massive schemas. Here is an example of the query rewriting process. {newshaded}User Query $Q_{0}$: Which semester the master and the bachelor both got enrolled in?
@@ 3.2 Modular Step Design
[2] Rewrite $Q_{2}$: In a database with enrollment_records, how to find semesters where both master and bachelor students enrolled? Group by semester and filter for overlapping enrollments.
Step two: isolate irrelevant schema information. While multi-round retrieval substantially enhances the recall of critical schema elements, embedding-based similarity comparisons are prone to introducing additional semantically proximate but irrelevant noise (Error 2). To mitigate this challenge, we introduce a filtering mechanism designed to prune redundant or irrelevant schema elements. Although we prioritize target database localization in multi-database settings, which challenges non-technical users who lack expert database knowledge, the framework remains effective in filtering out noise in single-database settings, which can serve as a subsequent operation after localization. To further improve performance in single-database settings, we propose two optimization strategies: Random Preservation with Exponential Decay and Post-Retrieval, detailed in the Appendix D. We now focus on the multi-database setting.
Once the retrieved results $Z$ contain schemas from irrelevant databases, the next step is to precisely locate the target database $D_{t}$ while filtering out irrelevant ones. To achieve this, the framework initially groups all schemas by their respective databases, enabling subsequent processing to treat each database as a cohesive unit. The framework then compares the relevance of each candidate database $D_{i}$ by evaluating how well its associated schemas satisfy the user’s query intent and then ranks these databases accordingly. The database exhibiting the highest relevance, $D_{t}$, is then designated as the target database, concurrently with the suppression of schema noise originating from irrelevant databases.
@@
where M denotes the LLMs used for analysis. By isolates unrelated schemas information, this step enables subsequent processes to concentrate computational resources on the most appropriate database, improving schema linking performance while achieving cost efficiency.
Step three: extract schemas for SQL generation. To mitigate schema misuse during SQL generation (Error 3 and Error 4), it is imperative to precisely parse the schema and identify the most relevant tables and columns. This procedure emulates the laborious manual process of schema extraction, yet it is fully automated by leveraging the intrinsic knowledge and reasoning capabilities of LLMs. Crucially, this approach scales schema linking to complex, redundant, and large-scale databases through the integration of advanced reasoning-enhanced prompting techniques, including multi-agent debate and chain-of-thought reasoning.
Specifically, from the filtered database schema $S_{\hat{u}}$ derived in the preceding steps, the objective is to identify a salient subset $S_{\hat{u}}^{\prime}$ comprising the most relevant tables $T_{i}^{\hat{u}}$ and columns $C_{i}^{\hat{u}}$. This selection is predicated on their alignment with the user query, thereby ensuring the resulting schema is both precise and comprehensively representative.
@@ 3.3 Component Implementation Optimization
Drawing upon the modular definitions outlined in Section 3.2, we introduce two distinct strategies to implement the core components of each step, as depicted by the dashed boxes in Figure 1. This modular framework enables flexible combinations based on specific query scenarios, allowing for optimized trade-offs between computational cost and processing effectiveness.
The first strategy is the Single-Prompt Pipeline, which executes each step through a single process-fixed LLM call. This design offers a low-latency streamlined solution, making it ideal for real-time database queries. A detailed exposition of this strategy is provided in Appendix C. Conversely, this section will primarily focus on the Multi-agent Collaboration strategy. This approach prioritizes accuracy and offers robust capabilities for tackling complex query tasks in real-world environments.
Align Semantics by Query Rewriting. Inspired by the reflective capabilities of LLMs demonstrated by Shinn et al., we introduce a semantic-enhanced retrieval approach based on retrieval feedback to achieve precise alignment with the ground-truth schema. Specifically, the Schema Auditor initiates by mapping the user queries into structured triplets (entities, attributes, and constraints). Next, it scrutinizes the retrieval results to infer missing schemas that may critical for accurate SQL generation (e.g., tables or fields for SELECT, JOIN, or WHERE clauses). This process culminates in an audit report that summarizes the parsed query, the inferred missing schemas, and the corresponding confidence levels. Subsequently, the Query Rewriter Agent leverages the comprehensive report to enhance the original query by disambiguating ambiguous expressions, supplementing semantic information for missing elements, and transforming the query into a template format optimized for text embedding models, ultimately improving retrieval recall for the ground-truth schema.
@@
Reduce Noise by Response Filtering. When multiple candidate schemas exhibit minimal semantic differentiation, achieving consensus through multi-agent debate can significantly mitigate the risk of confusion. Inspired by this insight, we meticulously designed a multi-agent debate model comprising two distinct LLM agents: Data Analyst and Database Expert. Specifically, the Data Analyst evaluates the alignment between each database and the user query based on domain relevance and schema coverage completeness, then ranking them through corresponding comprehensive assessment. The highest ranked database is then selected from all candidates, with its schema and query context provided to the Database Expert. Subsequently, the Database Expert rigorously evaluates whether its provided database schema can satisfy the query requirements, validating the selection’s appropriateness through pseudo-SQL generation and determining whether to retain it. The debate follows a one-by-one strategy, i.e., starting with the data analyst, after which the two roles present their perspectives in turn. The debate ends when a predefined number of rounds is reached, and then a terminator outputs the consensus database as a final result.
@@
Identify Tables and Columns by Schema Parsing. To enhance schema linking capabilities in complex scenarios, we meticulously designed a Multi-Agent Debate framework comprising two distinct LLM agents: Schema Parser and Data Scientist. Specifically, the Schema Parser extracts potentially required schema elements across three dimensions — tables, fields, and relationships — conducting reviews to prevent omissions. Extraction results from multiple Schema Parsers are then aggregated and submitted to the Data Scientist, who subsequently verifies all results, identifying any omissions or errors. The debate process follows a Simultaneous-Talk-with-Summarizer strategy, wherein multiple peer Schema Parsers engage in concurrent deliberation during each round, with final outcomes evaluated by the authoritative Data Scientist. Multi-role participation enhances the recall of tables and columns required for SQL generation, with diverse answers complementing each other to reduce the randomness of single-prompt outputs.